# Learning to Diagnose and Correct Moral Errors: Towards Enhancing Moral Sensitivity in Large Language Models

### v1

@@ Abstract
Bocheng Chen1111Equal contribution.    Han Zi2111Equal contribution.    Xi Chen3    Xitong Zhang4 Kristen Johnson4        Guangliang Liu4 1University of Mississippi  2Northeastern University  3Nanyang Technological University  4Michigan State University bchen5@olemiss.edu   zi.h@northeastern.edu   zoexi.chen@ntu.edu.sg {zhangxit,kristenj,liuguan5}@msu.edu
@@ 1 Introduction
According to  graff2024moral, moral sensitivity refers to the ability to recognize morally relevant features arising from moral situations and how those features are morally relevant. For example, associating males always with programmers and females always with nurses concerns gender discrimination, which is underpinned by the moral foundation of fairness 10.5555/3157382.3157584. Humans can effectively pick up these linguistic cues (e.g., male, female, occupations) and relate their associations to moral judgment, as an integral part of their moral competence moor2006nature. However, developing such moral sensitivity for machines has rather been onerous. Some doubt whether LLMs could actually achieve human-like moral sensitivity johnson2006computer; veliz2021moral; sparrow2021machines, while others argue for a more practical approach and abandon the pursuit of human-like moral sensitivity in machines anderson2011machine; anderson2014geneth; fossa2018artificial. From this perspective of practicalism, achieving moral sensitivity still remains an open question.
@@
The primary research focus of machine morality in the current LLM landscape aims to align LLMs with human moral values, through approaches such as reinforcement learning from human feedback bai2022training and moral self-correction ganguli2023capacity; liu2024intrinsic. But there have been studies arguing that these alignment techniques achieve only superficial moral response zhou2023lima; lee2024mechanistic; liu-etal-2025-discourse, without truly developing the LLMs’ moral sensitivity. Instead, moral sensitivity is sporadically observed in recent studies on LLMs’ moral reasoning ziems2022moral; jiang2025investigating; chen2025pragmatic; liu2025pragmatic. For example,  chen2025pragmatic proposes a multi-step pragmatic inference chain that prompts LLMs to detect implicit toxicity by recognizing features relevant to toxicity.
This study sees these sporadic findings as signs of LLMs’ potential in developing moral sensitivity. Drawn on the definition of moral sensitivity proposed by graff2024moral, it proposes pragmatic inference methods that train LLMs to diagnose and correct moral errors, thereby providing a pathway for improving LLMs’ moral sensitivity. The method design borrows insights from pragmatics in linguistics that studies how we understand “unsaid” meanings beyond literal meanings through a type of non-logical reasoning – pragmatic inference elder_pragmatic_2024. A related example is that the inference of moral wrongness of killing someone is not a result of a logical premise of one being killable or not, but is rather drawn on various contextual factors (e.g., the killer’s personality, social relationships) in a non-logical way. Moral reasoning is inherently pragmatic inferential liu2025pragmatic.
@@
Second, effective leverage of pragmatic information: LLMs often rely on superficial heuristics to seemingly make pragmatic inferences mccoy2019right; rogers-etal-2020-primer; shapira2024clever; sanchez2025metaphor. For example, in moral reasoning tasks, liu-etal-2025-diagnosing show that LLMs fail to leverage moral foundations that are provided and instead rely primarily on semantic cues to achieve generalization that is restricted by the type of task. Therefore, pragmatic inferences are not a simple plug-in that automatically enhance LLMs’ moral sensitivity, but rather need theoretically sound and experimentally verified design.
@@ 2 Related Works
As mentioned in Section 1, various approaches have been proposed to align LLMs with human moral values, including reinforcement learning from human feedback bai2022training and moral self-correction ganguli2023capacity. However, the alignment techniques have often been criticized for capturing only the mapping patterns (through statistical association) between a prompt and a moral judgment  zhou2023lima; liu2024intrinsic; lee2024mechanistic; liu-etal-2025-discourse; qi2024safety; lin2023unlocking; preniqi2024moralbert; tennant2025moral, while developing little sensitivity to what accounts for the moral judgment and how the judgment is made  reynolds2015recognition; cervantes2020artificial; sachdeva2025normative. In the meanwhile, although pragmatic inferences have been proven to be effective in detecting toxic language and social bias tasks  sap2020social; chen2025pragmatic, very few has examined whether it can be used for diagnosing and correcting moral errors and to what extent it is generalizable across different tasks.
More studies focus on evaluating the moral values and sociocultural norms embedded in LLMs scherrer2023evaluating; ramezani2023knowledge; adilazuarda2024towards; moore2024large. Since the Moral Foundation Theory (MFT) haidt2004intuitive; graham2013moral is widely used to characterize moral issues and understand human morality, several studies also investigate whether LLMs can recognize the MFs simmons2023moral; preniqi2024moralbert; abdulhai2024moral; zangari2025survey.  liu2025pragmatic is one of the first that incorporates MFT into pragmatic inference and successfully train LLMs to draw on them when making moral judgments. This study builds on their findings, but goes beyond their design of pragmatic inferences to dive into the development of LLMs’ moral sensitivity through diagnosing and correcting.
@@ 3.2 Motivation
Designing a method that facilitates LLMs to be morally sensitive needs to first address the two challenges that we listed in Section 1: (1) generalization across tasks: the method needs to develop LLMs’ sensitivity to different moral discourses that are varied by language use, semantic distributions, and explicitness, and (2) effective leverage of pragmatic information: the moral response produced by LLMs need to be a result of developed moral sensitivity, not that of heuristic pattern learning. On top of these challenges, there is a lack of established principles or guidelines on how to translate the one-task-one-way methods that previous studies have developed for toxic language, social bias, and security hazards, separately sap2020social; chen2025pragmatic, into one united yet flexibly adjustable approach.
Therefore, the first step we take to address the challenges is to consider moral foundations (MFs) as a generally applicable underpinning that connects different moral discourses. haidt2004intuitive and graham2013moral propose that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. They have been widely used in computational social science  johnson2018classification; wendell2021classifying; reinig2024survey, although rarely across different moral tasks.
However, MFs can only be a component involved in pragmatic inference, as shown in liu2025pragmatic. The second step towards a unified approach is thus to move beyond the semantic qualities of different moral discourses and analyze their pragmatic features liu-etal-2025-diagnosing. Specifically, we leave aside the different semantic distributions of different datasets, such as indirect social biases from the BBQ benchmark parrish-etal-2022-bbq, jailbreaks from the Jailbreakbench benchmark chao2024jailbreakbench and direct toxic language from the RealToxicityPrompts benchmark gehman2020realtoxicityprompts, to focus on the different levels of directness that their language choice creates. This step leads us to locate the variety of morality-related tasks to one common scale – pragmatic inference load – which is detailed in the design of learning2DiagCorr below.
@@ 3.3 The design of learning2DiagCorr
First, the inference load is determined by whether the moral discourse contains explicit linguistic cues that have a conventional indexicality with negative social qualities. Indexicality is a notion used across pragmatics, sociolinguistics, and anthropology. It refers to the connections that humans create between language use and social qualities silverstein_shifters_1976. The connections are developed over recurrent social experiences, with certain social qualities becoming more or less “fixed” to a language form ochs_culture_1988, e.g., the word “c*nt” indexing obscene, foolish, and unpleasant qualities. Processing these conventionalized indexicalities often does not require much cognitive effort, hence being light in terms of pragmatic inference load. In contrast, a heavier inference load applies to the implicit toxic language, social biases, and security threats that do not have overtly harmful expression and are creative and context-dependent. Their inferences often consist of conscious metapragmatic commentary – the evaluations of and explanations for one’s language use verschueren_notes_2000. An obvious example is that humans consistently ask themselves what others mean by saying/doing something. In the current method design, we bring such metapragmatic commentary, which could stay in the human mind, to texts for machines to learn.
@@
This study accordingly defines the different inference loads by their pathways: heavy load via metapragmatic commentary (top in Table 2) and light load via conventional indexicality (bottom in Table 2). The straightforward design of the light-load pragmatic inference is to enable LLMs to readily utilize explicit linguistic cues and prevent them from overthinking due to any unnecessary input of longer reasoning during the fine-tuning process. The heavy-load pragmatic inference, on the other hand, builds on the findings of liu2025pragmatic and develops five steps for moral diagnosis and correction.
@@
liu2025pragmatic finds that the inference of actions helps LLMs to clarify the target for moral judgment, the inference of action consequences brings about the models’ understanding of social norms, and moral foundations are effective in mediating the above two and the eventual moral judgment. We keep largely their basic design of action, consequence, and their connections to MFs with small changes in the current Steps 1 and 3, while adding new steps 4 and 5 for moral diagnosis and correction.
@@
Step 1 now includes “directly mentioned” and “implied” actions of the reply (the moral situation) – an idea inspired by speech act studies in pragmatics blum-kulka_requests_1984. It instructs the LLMs to take into consideration the prompt when deciding the actions performed by its reply. Step 3 replaces “which action associates with moral foundations” in liu2025pragmatic with “why the reply associates with moral foundations”. As stated in Section 3.1, our goal is to diagnose the reply, not the action. Steps 4 and 5 differ saliently from the steps in liu2025pragmatic. Step 4 affords the models to learn, not the moral evaluation, but the evaluative process of moral goodness and wrongness, which is manifested in the form of metapragmatic commentary on the links between action consequences, moral foundations, and moral judgments. Step 5 then teaches the models to utilize what they have learned in Step 1-4 to decide the editable part of the Reply and re-evaluate the revised reply.
@@
Notably, moral corrections do not necessarily contain a diagnosis. Models can still correct based on the learning of statistical heuristics liu-etal-2025-discourse. Our Step 5 intentionally requests the models to use the moral foundations violated by the immoral actions, identified in the diagnosis phase (Step 4), to carry out the correction, thereby explicitly linking diagnosis and correction through moral foundations. We will have further empirical evidence in Section 5 to demonstrate that learning2DiagCorr does not rely on heuristics for acquiring moral sensitivity.
@@ 4 Experiment
In this section, we introduce our experimental setting and results. To achieve generalization across tasks, we deliberately train the models on moral reasoning data MIC ziems2022moral and part of RealToxicPrompts gehman2020realtoxicityprompts, but test on BBQ (indirect social biases) and jailbreakbench as well as an unseen part of RealToxicPrompts that have been made different in data distribution. The experimental results demonstrate that (1) learning2DiagCorr achieves the best performance across the different tasks, and (2) learning2DiagCorr can generate high-quality diagnostic information that can even benefit off-the-shelf LLMs to correct moral errors through direct prompting.
@@ 4.1 Experimental Setting
Benchmarks. We employ three representative morality-relevant benchmarks: BBQ parrish-etal-2022-bbq, Jailbreakbench chao2024jailbreakbench, and RealToxicityPrompts benchmark gehman2020realtoxicityprompts to evaluate our learning2DiagCorr. We select these three benchmarks because represent tasks with diverse discourse characteristics, enabling a thorough evaluation of different pragmatic methods. For example, the indirect social biases in BBQ allow us to assess the effectiveness of our heavy-load pragmatic inference, while the direct toxic language in RealToxicityPrompts evaluates the performance of our light-load pragmatic inference. Jailbreaks from JailbreakBench combine both indirect and direct language, reflecting the characteristics of the other two benchmarks; for this dataset, we further analyze the benefits of integrating both heavy-load and light-load pragmatic inference methods.
Pragmatic Inference Fine-tuning. We leverage the popular moral reasoning dataset MIC ziems2022moral to fine-tune Llama-3.2-1Bhttps://huggingface.co/meta-llama/Llama-3.2-1B and Llama-3.2-3B§§§https://huggingface.co/meta-llama/Llama-3.2-3B base models, which serve as our backbone models for heavy-load pragmatic inference. The MIC benchmark provides annotations for moral foundations and revised replies, making it well-suited for heavy-load pragmatic inference. For light-load pragmatic inference, we fine-tune the same base models on the RealToxicityPrompts benchmark. As this benchmark does not include annotated revised replies, we design a prompting strategy to guide off-the-shelf LLMs in generating morally correct replies. Please refer to Appendix A.2 for more details about training pragmatic inference models.
@@
Direct Prompting (Direct): prompting LLMs to directly revise the reply and correct moral errors. Since direct prompting can only leverage the internal knowledge of LLMs, it tells how well the model can correct the reply itself, we consider it as a proxy of self-correction strategy liu2024intrinsic.
Chain-of-Thought (CoT): CoT prompting is a well-established strategy for improving performance on reasoning tasks. It similarly prompt the models to divide the task into several intermediate steps wei2023chainofthoughtpromptingelicitsreasoning. We treat it as a strong baseline against which our pragmatic inference methods are compared. Following the same data collection procedure used for learning2DiagCorr, we apply a standard CoT prompt to the same off-the-shelf LLMs and use the resulting responses as supervision to fine-tune the same base models. More details of the CoT prompts are in Appendix A.2.

### v2

@@ Abstract
Bocheng Chen1111Equal contribution.    Xi Chen2111Equal contribution.    Han Zi3    Haitao Mao4222This work was conducted at Michigan State University.    Zimo Qi5    Xitong Zhang6   Kristen Johnson6    Guangliang Liu6 1University of Mississippi    2Nanyang Technological University    3Northeastern University    4AWS AI Labs    5Johns Hopkins University    6Michigan State University bchen5@olemiss.edu   zi.h@northeastern.edu   zoexi.chen@ntu.edu.sg zqi15@jh.edu    {haitaoma,zhangxit,kristenj,liuguan5}@msu.edu
@@ 1 Introduction
According to  Graff (2024), moral sensitivity refers to the ability to recognize morally relevant features arising from moral situations and how those features are morally relevant. For example, associating males always with programmers and females always with nurses concerns gender discrimination, which is underpinned by the moral foundation of fairness Bolukbasi et al. (2016). Humans can effectively pick up these linguistic cues (e.g., male, female, occupations) and relate their associations to moral judgment, as an integral part of their moral competence Moor (2006). However, developing such moral sensitivity for machines has rather been onerous. Some doubt whether LLMs could actually achieve human-like moral sensitivity Johnson (2006); Véliz (2021); Sparrow (2021), while others argue for a more practical approach and abandon the pursuit of human-like moral sensitivity in machines Anderson and Anderson (2011, 2014); Fossa (2018). From this perspective of practicalism, achieving moral sensitivity still remains an open question.
@@
The primary research focus of machine morality in the current LLM landscape aims to align LLMs with human moral values, through approaches such as reinforcement learning from human feedback Bai et al. (2022) and moral self-correction Ganguli et al. (2023); Liu et al. (2024a). But there have been studies arguing that these alignment techniques achieve only superficial moral response Zhou et al. (2023); Lee et al. (2024); Liu et al. (2025c), without truly developing the LLMs’ moral sensitivity. Instead, moral sensitivity is sporadically observed in recent studies on LLMs’ moral reasoning Ziems et al. (2022); Jiang et al. (2025); Chen and Wang (2025); Liu et al. (2025a). For example,  Chen and Wang (2025) proposes a multi-step pragmatic inference chain that prompts LLMs to detect implicit toxicity by recognizing features relevant to toxicity.
This study sees these sporadic findings as signs of LLMs’ potential in developing moral sensitivity. Drawn on the definition of moral sensitivity proposed by Graff (2024), it proposes pragmatic inference methods that train LLMs to diagnose and correct moral errors, thereby providing a pathway for improving LLMs’ moral sensitivity. The method design borrows insights from pragmatics in linguistics that studies how we understand “unsaid” meanings beyond literal meanings through a type of non-logical reasoning – pragmatic inference Elder (2024). A related example is that the inference of moral wrongness of killing someone is not a result of a logical premise of one being killable or not, but is rather drawn on various contextual factors (e.g., the killer’s personality, social relationships) in a non-logical way. Moral reasoning is inherently pragmatic inferential Liu et al. (2025a).
@@
Second, effective leverage of pragmatic information: LLMs often rely on superficial heuristics to seemingly make pragmatic inferences McCoy et al. (2019); Rogers et al. (2020); Shapira et al. (2024); Sanchez-Bayona and Agerri (2025). For example, in moral reasoning tasks, Liu et al. (2025b) show that LLMs fail to leverage moral foundations that are provided and instead rely primarily on semantic cues to achieve generalization that is restricted by the type of task. Therefore, pragmatic inferences are not a simple plug-in that automatically enhance LLMs’ moral sensitivity, but rather need theoretically sound and experimentally verified design.
@@ 2 Related Works
As mentioned in Section 1, various approaches have been proposed to align LLMs with human moral values, including reinforcement learning from human feedback Bai et al. (2022) and moral self-correction Ganguli et al. (2023). However, the alignment techniques have often been criticized for capturing only the mapping patterns (through statistical association) between a prompt and a moral judgment  Zhou et al. (2023); Liu et al. (2024a); Lee et al. (2024); Liu et al. (2025c); Qi et al. (2024); Lin et al. (2023); Preniqi et al. (2024); Tennant et al. (2025), while developing little sensitivity to what accounts for the moral judgment and how the judgment is made  Reynolds and Miller (2015); Cervantes et al. (2020); Sachdeva and van Nuenen (2025). In the meanwhile, although pragmatic inferences have been proven to be effective in detecting toxic language and social bias tasks  Sap et al. (2020); Chen and Wang (2025), very few has examined whether it can be used for diagnosing and correcting moral errors and to what extent it is generalizable across different tasks.
More studies focus on evaluating the moral values and sociocultural norms embedded in LLMs Scherrer et al. (2023); Ramezani and Xu (2023); Adilazuarda et al. (2024); Moore et al. (2024). Since the Moral Foundation Theory (MFT) Haidt and Joseph (2004); Graham et al. (2013) is widely used to characterize moral issues and understand human morality, several studies also investigate whether LLMs can recognize the MFs Simmons (2023); Preniqi et al. (2024); Abdulhai et al. (2024); Zangari et al. (2025).  Liu et al. (2025a) is one of the first that incorporates MFT into pragmatic inference and successfully train LLMs to draw on them when making moral judgments. This study builds on their findings, but goes beyond their design of pragmatic inferences to dive into the development of LLMs’ moral sensitivity through diagnosing and correcting.
@@ 3.2 Motivation
Designing a method that facilitates LLMs to be morally sensitive needs to first address the two challenges that we listed in Section 1: (1) generalization across tasks: the method needs to develop LLMs’ sensitivity to different moral discourses that are varied by language use, semantic distributions, and explicitness, and (2) effective leverage of pragmatic information: the moral response produced by LLMs need to be a result of developed moral sensitivity, not that of heuristic pattern learning. On top of these challenges, there is a lack of established principles or guidelines on how to translate the one-task-one-way methods that previous studies have developed for toxic language, social bias, and security hazards, separately Sap et al. (2020); Chen and Wang (2025), into one united yet flexibly adjustable approach.
Therefore, the first step we take to address the challenges is to consider moral foundations (MFs) as a generally applicable underpinning that connects different moral discourses. Haidt and Joseph (2004) and Graham et al. (2013) propose that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. They have been widely used in computational social science  Johnson and Goldwasser (2018); Wendell and Tatalovich (2021); Reinig et al. (2024), although rarely across different moral tasks.
However, MFs can only be a component involved in pragmatic inference, as shown in Liu et al. (2025a). The second step towards a unified approach is thus to move beyond the semantic qualities of different moral discourses and analyze their pragmatic features Liu et al. (2025b). Specifically, we leave aside the different semantic distributions of different datasets, such as indirect social biases from the BBQ benchmark Parrish et al. (2022), jailbreaks from the Jailbreakbench benchmark Chao et al. (2024) and direct toxic language from the RealToxicityPrompts benchmark Gehman et al. (2020), to focus on the different levels of directness that their language choice creates. This step leads us to locate the variety of morality-related tasks to one common scale – pragmatic inference load – which is detailed in the design of learning2DiagCorr below.
@@ 3.3 The design of learning2DiagCorr
First, the inference load is determined by whether the moral discourse contains explicit linguistic cues that have a conventional indexicality with negative social qualities. Indexicality is a notion used across pragmatics, sociolinguistics, and anthropology. It refers to the connections that humans create between language use and social qualities Silverstein (1976). The connections are developed over recurrent social experiences, with certain social qualities becoming more or less “fixed” to a language form Ochs (1988), e.g., the word “c*nt” indexing obscene, foolish, and unpleasant qualities. Processing these conventionalized indexicalities often does not require much cognitive effort, hence being light in terms of pragmatic inference load. In contrast, a heavier inference load applies to the implicit toxic language, social biases, and security threats that do not have overtly harmful expression and are creative and context-dependent. Their inferences often consist of conscious metapragmatic commentary – the evaluations of and explanations for one’s language use Verschueren (2000). An obvious example is that humans consistently ask themselves what others mean by saying/doing something. In the current method design, we bring such metapragmatic commentary, which could stay in the human mind, to texts for machines to learn.
@@
This study accordingly defines the different inference loads by their pathways: heavy load via metapragmatic commentary (top in Table 2) and light load via conventional indexicality (bottom in Table 2). The straightforward design of the light-load pragmatic inference is to enable LLMs to readily utilize explicit linguistic cues and prevent them from overthinking due to any unnecessary input of longer reasoning during the fine-tuning process. The heavy-load pragmatic inference, on the other hand, builds on the findings of Liu et al. (2025a) and develops five steps for moral diagnosis and correction.
@@
Liu et al. (2025a) finds that the inference of actions helps LLMs to clarify the target for moral judgment, the inference of action consequences brings about the models’ understanding of social norms, and moral foundations are effective in mediating the above two and the eventual moral judgment. We keep largely their basic design of action, consequence, and their connections to MFs with small changes in the current Steps 1 and 3, while adding new steps 4 and 5 for moral diagnosis and correction.
@@
Step 1 now includes “directly mentioned” and “implied” actions of the reply (the moral situation) – an idea inspired by speech act studies in pragmatics Blum-Kulka and Olshtain (1984). It instructs the LLMs to take into consideration the prompt when deciding the actions performed by its reply. Step 3 replaces “which action associates with moral foundations” in Liu et al. (2025a) with “why the reply associates with moral foundations”. As stated in Section 3.1, our goal is to diagnose the reply, not the action. Steps 4 and 5 differ saliently from the steps in Liu et al. (2025a). Step 4 affords the models to learn, not the moral evaluation, but the evaluative process of moral goodness and wrongness, which is manifested in the form of metapragmatic commentary on the links between action consequences, moral foundations, and moral judgments. Step 5 then teaches the models to utilize what they have learned in Step 1-4 to decide the editable part of the Reply and re-evaluate the revised reply.
@@
Notably, moral corrections do not necessarily contain a diagnosis. Models can still correct based on the learning of statistical heuristics Liu et al. (2025c). Our Step 5 intentionally requests the models to use the moral foundations violated by the immoral actions, identified in the diagnosis phase (Step 4), to carry out the correction, thereby explicitly linking diagnosis and correction through moral foundations. We will have further empirical evidence in Section 5 to demonstrate that learning2DiagCorr does not rely on heuristics for acquiring moral sensitivity.
@@ 4 Experiment
In this section, we introduce our experimental setting and results. To achieve generalization across tasks, we deliberately train the models on moral reasoning data MIC Ziems et al. (2022) and part of RealToxicPrompts Gehman et al. (2020), but test on BBQ (indirect social biases) and jailbreakbench as well as an unseen part of RealToxicPrompts that have been made different in data distribution. The experimental results demonstrate that (1) learning2DiagCorr achieves the best performance across the different tasks, and (2) learning2DiagCorr can generate high-quality diagnostic information that can even benefit off-the-shelf LLMs to correct moral errors through direct prompting.
@@ 4.1 Experimental Setting
Benchmarks. We employ three representative morality-relevant benchmarks: BBQ Parrish et al. (2022), Jailbreakbench Chao et al. (2024), and RealToxicityPrompts benchmark Gehman et al. (2020) to evaluate our learning2DiagCorr. We select these three benchmarks because represent tasks with diverse discourse characteristics, enabling a thorough evaluation of different pragmatic methods. For example, the indirect social biases in BBQ allow us to assess the effectiveness of our heavy-load pragmatic inference, while the direct toxic language in RealToxicityPrompts evaluates the performance of our light-load pragmatic inference. Jailbreaks from JailbreakBench combine both indirect and direct language, reflecting the characteristics of the other two benchmarks; for this dataset, we further analyze the benefits of integrating both heavy-load and light-load pragmatic inference methods.
Pragmatic Inference Fine-tuning. We leverage the popular moral reasoning dataset MIC Ziems et al. (2022) to fine-tune Llama-3.2-1Bhttps://huggingface.co/meta-llama/Llama-3.2-1B and Llama-3.2-3B§§§https://huggingface.co/meta-llama/Llama-3.2-3B base models, which serve as our backbone models for heavy-load pragmatic inference. The MIC benchmark provides annotations for moral foundations and revised replies, making it well-suited for heavy-load pragmatic inference. For light-load pragmatic inference, we fine-tune the same base models on the RealToxicityPrompts benchmark. As this benchmark does not include annotated revised replies, we design a prompting strategy to guide off-the-shelf LLMs in generating morally correct replies. Please refer to Appendix A.2 for more details about training pragmatic inference models.
@@
Direct Prompting (Direct): prompting LLMs to directly revise the reply and correct moral errors. Since direct prompting can only leverage the internal knowledge of LLMs, it tells how well the model can correct the reply itself, we consider it as a proxy of self-correction strategy Liu et al. (2024a).
Chain-of-Thought (CoT): CoT prompting is a well-established strategy for improving performance on reasoning tasks. It similarly prompt the models to divide the task into several intermediate steps Wei et al. (2023). We treat it as a strong baseline against which our pragmatic inference methods are compared. Following the same data collection procedure used for learning2DiagCorr, we apply a standard CoT prompt to the same off-the-shelf LLMs and use the resulting responses as supervision to fine-tune the same base models. More details of the CoT prompts are in Appendix A.2.